The probability of getting at least one head when a coin is tossed twice is:
A$\tfrac14$
B$\tfrac34$
C$1$
D$\tfrac12$
Answer & Solution
Correct answer: B. $\tfrac34$
P(at least one head) = 1 − P(no head) = 1 − 1/4 = 3/4.
Keep practising Application of Integrals
Browse all Application of Integrals questions for AP Intermediate 2nd Year, or explore more Mathematics topics.
Related questions